OpenStudy (anonymous):

there are 3 boats A, B, C. boat B has 15 more rooms than boat A. boat C has 39 fewer than twice the rooms of boat A. there is a total of 388 rooms. how many room do each boat have?

OpenStudy (anonymous):

Try breaking the paragraph up. Each sentence corresponds to an equation. \[ \text{Let }r_1\text{ be the number of rooms in boat A, etc.} Then use the equations to solve for each r

OpenStudy (anonymous):

bleh. forgot to close my TeX.

OpenStudy (anonymous):

\[ r_1 \text{ is the number of rooms on boat A, etc. }\]

OpenStudy (anonymous):

So for the first sentence: "Boat B has 15 more rooms than boat A." What would the equation representing this statement look like?

OpenStudy (anonymous):

b+15=a

OpenStudy (anonymous):

Very nearly. From your equation which boat has the greater number of rooms? And does that coincide with the information presented in the statement?

OpenStudy (anonymous):

from what i under stand: A=? B=15+A C=2A-39

OpenStudy (anonymous):

Yes, that's it exactly. Now we know one more thing which is that there is a total of 388 rooms. So write the equation for that statement.

OpenStudy (anonymous):

The last equation would be ?

OpenStudy (anonymous):

A+B+C=388

OpenStudy (anonymous):

sorry had to step away real quick

OpenStudy (anonymous):

Right. So now take what you know about C and B and replace those variables with expressions that have A instead.

OpenStudy (anonymous):

A+15+A+2A-39=388

OpenStudy (anonymous):

Yep. Then solve for A.

OpenStudy (anonymous):

Once you know A, you can go back to the equations for B and C to find what those are.

OpenStudy (anonymous):

i got it from here thanks a lot!

OpenStudy (anonymous):

Certainly. Just remember that each sentence is telling you something. Typically it translates directly into an equation, and by using all your equations you can do substitution to solve for one of your variables. Then once you have one you can find the others.